# Meridional circulation molecular-weighted

### v1

@@ Abstract
Meridional circulation in stratified stellar/planetary interiors in the presence of stable molecular-weight gradients remains poorly understood, thereby affecting angular momentum transport in evolutionary models. We extend the downward control principle of atmospheric sciences to include compositional stratification. Using a linearized analysis we show that stable $\mu$-gradients slow down the penetration of circulation into the depths, emphasizing the importance of time-dependent solutions. However, additional effects such as horizontal turbulence or magnetic fields are needed to halt it completely. We also find limits demarcating linear and nonlinear regimes in terms of Schmidt and Rossby numbers. Nonlinear simulations exhibit compositional mixing due to meridional currents, enabling deeper penetration than otherwise. We propose slowly evolving and steady-state for the solar tachocline, and helioseismically observed heavy element abundances, acknowledging the absence of constraints on the radial variation of the Schmidt number. In the context of stellar evolution, differential rotation profiles of solar-type main-sequence stars may follow analytical solutions extending from Banik & Menou (2024), thereby aiding further probes into magneto/hydrodynamic instabilities and outward angular momentum transport.
@@ 1 Introduction
Figure 2: The process of penetration of zonal forcing (top-left panel) and corresponding meridional circulation (circular lines with arrows), under effects of pure rotation, thermal stratification, compositional stratification, and horizontal momentum transport via. eddy/molecular viscosity. The individual panels resemble f-planes similar to the computational domain shown later in Figure 3. (Top left) Gyroscopic pumping: Blue and white regions correspond to regions of opposing zonal velocity. The zonal velocity and meridional circulation penetrate all the way together (Showman et al., 2006). (Top right) Radiative penetration: The yellow isotherms are tilted by the adiabatic meridional circulation, thereby resisting the penetration, which eventually proceeds due to radiative diffusion (Spiegel & Zahn, 1992). (Bottom left) Compositional-thermal wind: The magenta lines of constant mean-molecular weight are tilted in the opposite direction by the penetrating flow, which additionally resists the flow making the penetration even slower, yet reaching the bottom at steady-state. (Bottom right) Horizontal turbulence/viscosity: Horizontal momentum transport by eddy viscosity or magnetic fields oppose the penetration. Red and blue regions represent fluid of relatively higher and lower temperatures. Dark grey and light grey regions represent higher and lower mean-molecular weight respectively. The green region is a technically approximate representation of the opposition offered by viscous effects.
@@ Acknowledgements
D.B. thanks Thaddeus D. Komacek and Marta L. Bryan for the extensive discussions and feedback on the paper. D.B. is grateful to the anonymous reviewer whose comments improved the manuscript significantly. D.B. is supported by the Dept. of Physics, University of Toronto. K.M. is supported by the National Science and Engineering Research Council of Canada. E.H.A acknowledges partial support from NSF grant PHY-2309135 and Simons Foundation grant (216179, LB) to the Kavli Institute for Theoretical Physics (KITP).

### v2

@@ Abstract
Meridional circulation in stratified stellar/planetary interiors in the presence of stable molecular-weight gradients remains poorly understood, thereby affecting angular momentum transport in evolutionary models. We extend the downward control principle of atmospheric sciences to include compositional stratification. Using a linearized analysis we show that stable $\mu$-gradients slow down the penetration of circulation into the depths, emphasizing the importance of time-dependent solutions. However, additional effects such as horizontal turbulence or magnetic fields are needed to halt it completely. We also find limits demarcating linear and nonlinear regimes in terms of Schmidt and Rossby numbers. Nonlinear simulations exhibit compositional mixing due to meridional currents, enabling deeper penetration than otherwise. We propose slowly evolving and steady-state scenarios for the solar tachocline e and helioseismically observed heavy element abundances, while acknowledging the absence of constraints on the radial variation of the Schmidt number. In the context of stellar evolution, differential rotation profiles of solar-type main-sequence stars may follow analytical solutions extending from Banik & Menou (2024), thereby aiding further probes into magneto/hydrodynamic instabilities and outward angular momentum transport.
@@ 1 Introduction
Figure 2: The process of penetration of zonal forcing (top-left panel) and corresponding meridional circulation (circular lines with arrows) under effects of pure rotation, thermal stratification, compositional stratification, and horizontal momentum transport via. eddy/molecular viscosity. The individual panels resemble f-planes similar to the computational domain shown later in Figure 3. (Top left) Gyroscopic pumping: Blue and white regions correspond to regions of opposing zonal velocity. The zonal velocity and meridional circulation penetrate together all the way to the bottom of the domain following the Taylor Proudman theorem over a dynamical timescale (Showman et al., 2006). (Top right) Radiative penetration: The yellow isotherms from a stable thermal stratification are tilted by the adiabatic meridional circulation, thereby resisting the penetration, which eventually proceeds slowly due to radiative diffusion over the Eddington-Sweet timescale (Spiegel & Zahn, 1992). (Bottom left) Compositional-thermal wind: The magenta lines of constant mean-molecular weight from compositional stratification are tilted in the opposite direction by the meridional flow, additionally resisting it and making the penetration even slower, yet reaching the bottom at steady-state. (Bottom right) Horizontal turbulence/viscosity: Horizontal momentum transport by eddy viscosity or magnetic fields oppose the penetration, thereby halting it at a finite depth and rendering a faster steady-state (Spiegel & Zahn, 1992; Gough & McIntyre, 1998). Red and blue regions represent fluid of relatively higher and lower temperatures, and dark grey and light grey regions represent higher and lower mean-molecular weight, respectively. The green region is an approximate representation of the opposition offered by viscous or magnetic effects.
@@ Acknowledgements
D.B. thanks Thaddeus D. Komacek and Marta L. Bryan for the extensive discussions and feedback on the paper. D.B. is grateful to the anonymous reviewer whose comments improved the manuscript significantly. D.B. is supported by the Dept. of Physics, University of Toronto. D. B. also thanks Nicholas Brummell for related discussions on the topic. K.M. is supported by the National Science and Engineering Research Council of Canada. E.H.A acknowledges partial support from NSF grant PHY-2309135 and Simons Foundation grant (216179, LB) to the Kavli Institute for Theoretical Physics (KITP).